The Vinson Corporation has earnings of $979,000 with 340,000 shares outstanding. Its P/E ratio is 18. The firm is holding $460,000 of funds to invest or pay out in dividends. If the funds are retained, the aftertax return on investment will be 20 percent, and this will add to present earnings. The 20 percent is the normal return anticipated for the corporation, and the P/E ratio would remain unchanged. If the funds are paid out in the form of dividends, the P/E ratio will increase by 10 percent because the stockholders in this corporation have a preference for dividends over retained earnings.


a. Compute the price of the stock under the two plans. (Do not round intermediate calculations and round your answers to 2 decimal places.)
  

Retention Plan-

Payout Plan-

b. Which plan will maximize the market value of the stock?
  

Retention plan
Payout plan
Answer #1

Solution :

Earning = $979,000

Number of share = 340,000

Earning per share = $979,000 / 340,000 = 2.879412

P/E = Price per share / Earning per share = 18

Current price based on P/E ratio = P/E * EPS = 18 * 2.879412 = 51.8294

Now company has 460,000 of funds that can be used as dividend or be retained

So option 1 : When fund is retained

It is given that fund will generate 20% on 460,000 so extra earning = 20%* 460,000 = 92,000

This amount will be added to current earning

So new earning = 979,000 + 92,000 = 1,071,000

Number of share = 340,000

Earning per share = 1,071,000 /  340,000 = 3.15

P/E will remain same at 18

So price = P/E * EPS = 18 * 3.15 = $56.70

Option 2: When 460,000 is given as dividend

P/E will increase by 10% so

New P/E = 18* (1+10%) = 19.8

EPS will remain same at 2.879412 as there is no change in earnings

Share price =  P/E * EPS = 19.8 * 2.879412 = $57.01

Part B )

Dividend plan will maximize the market value of stock as it gives value of $57.01 as compared to retention plan which gives value of $56.7